OpenStudy (clamin):

Please help! Find S_n for each series described. You will need to determine if the series is arithmetic or geometric. 1. 160 + 80 + 40 + ....., n=6 2. d= -2/3, n= 16, a_n= 44

OpenStudy (anonymous):

looks like they are dividing by two each time for the first one

OpenStudy (clamin):

How do i solve it? @satellite73

OpenStudy (anonymous):

This should be a geometric progression since the difference in two terms is not constant You would need to find the (r) ratio in which the terms decrease, this can be done by dividing the second term by the first term r = 1/2 the formula for the sum of geometric progression \[S_{n} = \frac{a(1-r^{n})}{1-r}\]

OpenStudy (clamin):

Which is it for @thushananth01? Question 1 or 2?

OpenStudy (anonymous):

1

OpenStudy (clamin):

How do i solve the 2nd? @thushananth01

OpenStudy (anonymous):

Use the formula for arithmetic progression \[S_n = \frac{n}{2}(2a + (n-1)d)\]

OpenStudy (clamin):

What do i substitute with 2a?? @thushananth01

OpenStudy (anonymous):

That should be a_n

OpenStudy (clamin):

Ok